# Inference-Aware Prompt Optimization for Aligning Black-Box Large Language Models

### v1

@@ Acknowledgments
The proof follows directly (Karnin, Koren, and Somekh 2013) Lemma 4.3. The only thing to recognize is that:
$\mathbb{E}[N_{r}]$ $=\sum_{a_{i}\in\mathcal{A}^{(r)}_{c}}\Pr\!\big{[}\hat{Q}^{\alpha,(r)}_{c,a_{1}}<\hat{Q}^{\alpha,(r)}_{c,a_{i}}\big{]}$
$\leq\sum_{a_{i}\in\mathcal{A}^{(r)}_{c}}\exp\!\Big{(}-\Delta_{c,a_{i}}^{2}\cdot\tfrac{2^{r}T}{8|\mathcal{C}||\mathcal{A}|\bar{N}_{i}\log_{2}|\mathcal{A}|}\Big{)}$
@@
$3|\mathcal{C}|R\exp\!\Big{(}-\tfrac{T}{min(2|\mathcal{P}|H_{1},8|\mathcal{C}|H_{2})R}\Big{)}$
@@
The Inference-Agnostic prompt-optimization policy remains optimal under linear transformation of $R_{x}^{\textsc{IA}}(c,a)$, that is, $kR_{x}^{\textsc{IA}}(c,a),k>0$ and an optimal policy can be recovered trivially from $Q$-function under affine transformation:
$Q^{AF}(c,a):=\mathbb{E}_{x\sim\mathcal{X}}\bigl{[}aR_{x}^{IA}(c,a)+b\bigr{]}=kQ^{IA}(c,a)+b.\quad$
@@
Follows directly from Jensen’s inequality. ∎
@@
A pull of $N\leq N_{\max}$ for prompt $p$ on example $x$ generates i.i.d. Bernoulli outcomes $\{c_{i}\}_{i=1}^{N}$ where $\Pr(c_{i}=1)=q_{p}(x)$, and each completion incurs a per-sample cost $k_{p}$. The result is an array $\bigl{[}\;c_{i},\;k_{p}\;\bigr{]}_{i=1}^{N}$.
Majority vote (MV) sets $M=1$ if $\sum_{i}c_{i}>N/2$, $M=0$ if $\sum_{i}c_{i}<N/2$, and assigns $M=0.5$ (by fair coin) in the case of a tie ($N$ even, $\sum_{i}c_{i}=N/2$).
The utility for cost for context $c\in\{\text{low},\text{mid},\text{high}\}$ is computed as
$u_{c}=w_{1}M+w_{2}(c)\sum_{i=1}^{N}k_{p},$
where $w_{1}=1$ and $w_{2}(c)\in\{0,\ -0.2,\ -1.0\}$ depending on the cost tier.
@@
We model $P=32$ prompts, each paired with $|\mathcal{X}|=512$ queries and $K=2$ positive objectives. For every $(p,x)$, there are $M$ categorical outcomes, each represented by a vector $o_{j}\in\mathbb{R}^{K}$. A pull of $N\leq N_{\max}(=32)$ for prompt $p$ on query $x$ generates $N$ i.i.d. outcome vectors, resulting in rows $o_{i,1},\,o_{i,2},\,k_{p}\;]$, where $k_{p}$ denotes the per-completion cost for prompt $p$.
Given a context $c$ with weights $w=(w_{1},w_{2},w_{\text{cost}})$, where $w_{1}+w_{2}=1$ and $w_{\text{cost}}\leq 0$, the Best-of-$N$ utility is defined as
$u_{c}=\max_{1\leq i\leq N}\left(w_{1}o_{i,1}+w_{2}o_{i,2}\right)+w_{\text{cost}}\,N\,k_{p}.$
To construct the environment, outcome vectors are sampled from $\{-4,\ldots,4\}^{2}$. We instantiate two prompt archetypes: HMLV (high mean, low variance; excels at N=1N{=}1) and LMHV (lower mean, high variance; benefits from larger $N$), each specializing in one objective. For each $(p,x)$, we add small per-query noise to the categorical outcome probabilities, introduce a mild train-to-test shift by perturbing these probabilities, sample per-prompt costs $k_{p}\in[0.02,\,0.1]$, and draw context weights from a grid satisfying $w_{1}+w_{2}=1$ with $w_{\text{cost}}\in\{-0.1,\ -0.5,\ -1.0\}$.
@@
    For each problem, retain the global top-4 answers and group all other answers into a single other bucket ($C=5$ categories in total).
@@
This yields a categorical environment (analogous to the Synthetic-Categorical setting) with $P=25$, $N_{\max}=32$, a uniform context prior $c\in\{\text{low},\,\text{mid},\,\text{high}\}$, and cost coefficients {0, $\{0,\,0.2,\,1$.0\}. Utility is evaluated via majority vote.
@@
We randomly sample 1,5001{,}500 multiple-choice questions from the CommonsenseQA corpus [^3] https://huggingface.co/datasets/tau/commonsense˙qa, and author $48$ prompt templates using ChatGPT-o3. For each $(\text{prompt},\,\text{question})$ pair, we query Llama-3.3-70B-Instruct at temperature $T=1.1$, collecting $128$ JSON-constrained answers (one of “Option A”–“Option E”). Each prompt is assigned a constant cost $k_{p}=0.01$.
The resulting data is used to construct a categorical environment (in analogy to the Synthetic-Categorical setting) with $P=48$, $N_{\max}=32$, a uniform context prior, and cost coefficients {0, $\{0,\,0.2,\,1$.0\}.
@@
The two reward scores are then binned on a $0.5$-spaced grid, producing a categorical distribution per $(\text{prompt},\,\text{query})$; per-prompt costs are computed as the average token length. This data defines a categorical environment with $P=20$, $N_{\max}=32$, a uniform context prior over weight triples $(w_{h},\,w_{s},\,w_{\text{cost}})$ with $w_{h}+w_{s}=1$ and $w_{\text{cost}}\in\{-0.1,\,-0.5,\,-1.0\}$.
@@
This data defines a categorical environment with $P=20$, $N_{\max}=32$, and a uniform context prior over weight triples $(w_{h},\,w_{s},\,w_{\text{cost}})$ where $w_{h}+w_{s}=1$ and $w_{\text{cost}}\in\{-0.1,\,-0.5,\,-1.0\}$.
@@
For the screening variant, we fixed $K=4$ candidates after screening and swept the burn-in fraction $\rho\in\{0.05,0.10,0.20,0.30,0.40\}$, which allocates a $\rho$-portion of the budget to obtain initial estimates before trimming. Parameter sweep protocol matched the baselines. We selected $\rho=0.20$ for reporting, as it achieved the best overall performance while remaining robust across datasets and inference regimes 2.
@@
We tuned the exploration constant over $c\in\{0.1,0.5,1.0,2.0,4.0,8.0\}$ under the same budgets and using 20% of the data per environment; identical seeds across settings; 10,00010{,}000 test contexts). The agent ranks arms by the standard UCB index
@@
where $\hat{\mu}_{i}(t)$ is the empirical mean utility of arm $i$, $n_{i}(t)$ its pull count, and $t$ the total pulls. We selected $c=0.1$ for reporting, as it achieved the best overall performance while remaining robust across datasets and inference regimes 3.
@@
We swept $\varepsilon\in\{0.50,0.75,0.80,0.85,0.90,0.95\}$ separately for each dataset and inference regime (MV, BoN). For every $\varepsilon$, agents were trained under budgets $T\in\{3\text{K},5\text{K},10\text{K},20\text{K},30\text{K},40\text{K}\}$, using 20% of the data per environment with deterministic reseeding; evaluation used 10,00010{,}000 test contexts per environment. We selected $\epsilon=0.15$ for reporting, as it achieved the best overall performance while remaining robust across datasets and inference regimes 4.
@@
All the results are shown in Figs 56789,  10. Across all six datasets, we observe that PSST and the Top-K screening heuristic consistently outperform competing methods across most budget settings, with statistical significance.

### v2

@@ Acknowledgments
This follows by adapting Lemma 4.3 of Karnin et al. (2013) to our setting (using the same indexing convention as above). In particular, the key intermediate step is that
$\mathbb{E}[N_{r}]$ $\leq\sum_{a^{c,i}\in\mathcal{A}_{c}^{(r)}}\Pr\!\big[\hat{Q}^{\alpha,(r)}(c,a^{c,1})<\hat{Q}^{\alpha,(r)}(c,a^{c,i})\big]$
$\leq\sum_{a^{c,i}\in\mathcal{A}_{c}^{(r)}}\exp\!\Bigl(-\Delta_{c,a^{c,i}}^{2}\cdot\tfrac{2^{r}T}{8|\mathcal{C}|\,|\mathcal{A}|\,\bar{N}_{i}\,\log_{2}|\mathcal{A}|}\Bigr),$
and proceeding exactly as in Karnin et al. (2013) yields the stated bound (with complexity term $H_{2}^{c}$). ∎
@@
$3|\mathcal{C}|R\exp\!\Big(-\tfrac{T}{\min(2|\mathcal{P}|H_{1},8|\mathcal{C}|H_{2})R}\Big)$
@@
The Inference-Agnostic prompt-optimization policy remains optimal under linear transformation of $R_{x}^{\textsc{IA}}(c,a)$, that is, $\sigma R_{x}^{\textsc{IA}}(c,a),\sigma>0$ and an optimal policy can be recovered trivially from $Q$-function under affine transformation:
$Q^{AF}(c,a):=\mathbb{E}_{x\sim\mathcal{X}}\bigl[\sigma R_{x}^{\textsc{IA}}(c,a)+\mu\bigr]=\sigma Q^{\textsc{IA}}(c,a)+\mu.\quad$
@@
By linearity of expectation,
$Q^{\prime}(c,a)$ $=\mathbb{E}_{x}\!\left[\sigma R_{x}^{\textsc{IA}}(c,a)+\mu\right]$
$=\sigma\,\mathbb{E}_{x}\!\left[R_{x}^{\textsc{IA}}(c,a)\right]+\mu$
$=\sigma Q^{\textsc{IA}}(c,a)+\mu.$
Since $\sigma>0$, for any two arms $a,b$ we have
$Q^{\prime}(c,a)\geq Q^{\prime}(c,b)$ $\iff\sigma Q^{\textsc{IA}}(c,a)+\mu\geq\sigma Q^{\textsc{IA}}(c,b)+\mu$
$\iff Q^{\textsc{IA}}(c,a)\geq Q^{\textsc{IA}}(c,b).$
Therefore the ordering of arms is preserved and the $\arg\max$ set is identical. ∎
@@
A pull of $N\leq N_{\max}$ for prompt $p$ on query $x$ generates i.i.d. Bernoulli outcomes $\{z_{i}\}_{i=1}^{N}$ where $\Pr(z_{i}=1)=q_{p}(x)$, and each completion incurs a per-completion cost $k_{p}$. The result is an array $\bigl[\;z_{i},\;k_{p}\;\bigr]_{i=1}^{N}$.
Majority Voting (MV) sets $M=1$ if $\sum_{i}z_{i}>N/2$, $M=0$ if $\sum_{i}z_{i}<N/2$, and assigns $M=0.5$ (by fair coin) in the case of a tie ($N$ even, $\sum_{i}z_{i}=N/2$).
The cost-adjusted utility for context $c\in\{\text{low},\text{mid},\text{high}\}$ is computed as
$u_{c}=w_{1}M-w_{2}(c)\sum_{i=1}^{N}k_{p},$
where $w_{1}=1$ and $w_{2}(c)\in\{0,\ 0.2,\ 1.0\}$ depending on the cost tier.
@@
We model $|\mathcal{P}|=32$ prompts, each paired with $|\mathcal{X}|=512$ queries and $K=2$ positive objectives. For every $(p,x)$, there are $M$ categorical outcomes, each represented by a vector $o_{j}\in\mathbb{R}^{K}$. A pull of $N\leq N_{\max}$ (with $N_{\max}=32$) for prompt $p$ on query $x$ generates $N$ i.i.d. outcome vectors, resulting in rows $o_{i,1},\,o_{i,2},\,k_{p}\;]$, where $k_{p}$ denotes the per-completion cost for prompt $p$.
Given a context $c$ with weights $w=(w_{1},w_{2},w_{\text{cost}})$, where $w_{1}+w_{2}=1$ and $w_{\text{cost}}\geq 0$, the Best-of-$N$ (BoN) utility is defined as
$u_{c}=\max_{1\leq i\leq N}\left(w_{1}o_{i,1}+w_{2}o_{i,2}\right)-w_{\text{cost}}\,N\,k_{p}.$
To construct the environment, outcome vectors are sampled from $\{-4,\ldots,4\}^{2}$. We instantiate two prompt archetypes: HMLV (high mean, low variance; excels at N=1N{=}1) and LMHV (lower mean, high variance; benefits from larger $N$), each specializing in one objective. For each $(p,x)$, we add small per-query noise to the categorical outcome probabilities, introduce a mild train-to-test shift by perturbing these probabilities, sample per-prompt costs $k_{p}\in[0.02,\,0.1]$, and draw context weights from a grid satisfying $w_{1}+w_{2}=1$ with $w_{\text{cost}}\in\{0.1,\ 0.5,\ 1.0\}$.
@@
    For each problem, retain the global top-4 answers and group all other answers into a single other bucket (five categories in total).
@@
This yields a categorical environment (analogous to the Synthetic-Categorical setting) with $|\mathcal{P}|=25$, $N_{\max}=32$, a uniform context prior $c\in\{\text{low},\,\text{mid},\,\text{high}\}$, and cost coefficients {0, $\{0,\,0.2,\,1$.0\}. Utility is evaluated via MV.
@@
We randomly sample 1,5001{,}500 multiple-choice questions from the CommonsenseQA corpus [^3] https://huggingface.co/datasets/tau/commonsense˙qa, and author $48$ prompt templates using ChatGPT-o3. For each $(\text{prompt},\,\text{question})$ pair, we query Llama-3.3-70B-Instruct at temperature $T=1.1$, collecting $128$ JSON-constrained answers (one of “Option A”–“Option E”). Each prompt is assigned a constant cost $k_{p}=0.01$ as we use JSON mode, and all completion has equal length.
The resulting data is used to construct a categorical environment (in analogy to the Synthetic-Categorical setting) with $|\mathcal{P}|=48$, $N_{\max}=32$, a uniform context prior, and cost coefficients {0, $\{0,\,0.2,\,1$.0\}.
@@
The two reward scores are then binned on a $0.5$-spaced grid, producing a categorical distribution per $(\text{prompt},\,\text{query})$; per-prompt costs are computed as the average token length. This data defines a categorical environment with $|\mathcal{P}|=20$, $N_{\max}=32$, and a uniform context prior over weight triples $(w_{\text{help}},\,w_{\text{harm}},\,w_{\text{cost}})$ with $w_{\text{help}}+w_{\text{harm}}=1$ and $w_{\text{cost}}\in\{0.1,\,0.5,\,1.0\}$.
@@
This data defines a categorical environment with $|\mathcal{P}|=20$, $N_{\max}=32$, and a uniform context prior over weight triples $(w_{\text{pref}},\,w_{\text{faith}},\,w_{\text{cost}})$ where $w_{\text{pref}}+w_{\text{faith}}=1$ and $w_{\text{cost}}\in\{0.1,\,0.5,\,1.0\}$.
@@
For the screening variant, we fixed $K=4$ candidates after screening and swept the burn-in fraction $\rho\in\{0.05,0.10,0.20,0.30,0.40\}$, which allocates a $\rho$-fraction of the budget to obtain initial estimates before trimming. The parameter-sweep protocol matched the baselines. We selected $\rho=0.20$ for reporting, as it achieved the best overall performance while remaining robust across datasets and inference regimes (Table 2).
@@
We tuned the exploration constant over $c\in\{0.1,0.5,1.0,2.0,4.0,8.0\}$ under the same budgets, using 20% of the data per environment with identical seeds across settings, and 10,00010{,}000 test contexts. The agent ranks arms by the standard UCB index
@@
where $\hat{\mu}_{i}(t)$ is the empirical mean utility of arm $i$, $n_{i}(t)$ its pull count, and $t$ the total pulls. We selected $c=0.1$ for reporting, as it achieved the best overall performance while remaining robust across datasets and inference regimes (Table 3).
@@
We swept the greedy probability $1-\epsilon\in\{0.50,0.75,0.80,0.85,0.90,0.95\}$ separately for each dataset and inference regime (MV, BoN). For every $\epsilon$, agents were trained under budgets $T\in\{3\text{K},5\text{K},10\text{K},20\text{K},30\text{K},40\text{K}\}$, using 20% of the data per environment with deterministic reseeding; evaluation used 10,00010{,}000 test contexts per environment. We selected $\epsilon=0.15$ for reporting, as it achieved the best overall performance while remaining robust across datasets and inference regimes (Table 4).
@@
All the results are shown in Figs 678910,  11. Across all six datasets, we observe that PSST and the Top-K screening heuristic consistently outperform competing methods across most budget settings, with statistical significance.